OpenStudy (compassionate):

TUTORIAL: Taking request for a new tutorial. Most requested field or subject/topic wins. Go.

OpenStudy (compassionate):

I suggest a topic on integrals.

OpenStudy (anonymous):

Me too

OpenStudy (anonymous):

Integration by Parts?

mathslover (mathslover):

integration

OpenStudy (compassionate):

By parts anyone?

OpenStudy (anonymous):

I would love to learn diferential integrals......... cause I have no idea about that. and have seen some other guys doing that

OpenStudy (unklerhaukus):

From the product rule \[\frac{\text d\left(uv\right)}{\text dx} = \frac{\text du}{\text dx}v+u\frac{\text d v}{\text dx}\]\[ \frac{\text du}{\text dx}v+u\frac{\text d v}{\text dx}=\frac{\text d\left(uv\right)}{\text dx} \]\[ \frac{\text du}{\text dx}v=\frac{\text d\left(uv\right)}{\text dx} -u\frac{\text d v}{\text dx}\]\[ \int v\frac{\text du}{\text dx}\text dx=\int\frac{\text d\left(uv\right)}{\text dx}\text dx -u\int\frac{\text d v}{\text dx}\text dx\]\[ \int{v\text du}=\left. {uv}\right| -u\int {\text d v}\]Integration by Parts

OpenStudy (anonymous):

you could probably expand on this by discussing the strategies used in integrating the different combinations of functions. i.e. integrat(xlnx) and so on..